Consider the following AS-AD model. Please note that I have given you the AD curve explicitly and simplified many of the expressions you saw in lecture and previous HW (i.e. no z, I already solved for the relationship between unemployment and output, etc.).

Aggregate demand: Y = 200 − 10P

Wage setting relationship: w = P e (1 − u)

Price setting relationship: P = 1.1w

Output and the unemployment rate: u = 1 − Y 110

1. Derive the AS curve.
Answer #1

Answer 1

AS curve shows the quantity of output(GDP) that will be produced and supplied at different price level. Higher the price level more output will be supplied. This means that there is a positive relationship between Y(output or GDP) and P(Price level).

From Wage setting relationship we have w = Pe(1 − u) and from Output and the unemployment rate we have : u = 1 − Y/110

From above equations we have :

w = Pe(1 − u) = Pe(1 − (1 − Y/110)) = Pe(Y/110)

Putting this in price setting relationship we get :

P = 1.1w = 1.1Pe(Y/110)

=> Y = 100P/Pe -------------------------------------Aggregate supply(AS curve equation)

(Note that P is directly related to Y as discussed above)